# Multimodal Cross-Domain Few-Shot Learning for Egocentric Action Recognition

### v1

@@ Abstract
We address a novel cross-domain few-shot learning task (CD-FSL) with multimodal input and unlabeled target data for egocentric action recognition. This paper simultaneously tackles two critical challenges associated with egocentric action recognition in CD-FSL settings: (1) the extreme domain gap in egocentric videos (e.g., daily life vs. industrial domain) and (2) the computational cost for real-world applications. We propose MM-CDFSL, a domain-adaptive and computationally efficient approach designed to enhance adaptability to the target domain and improve inference speed. To address the first challenge, we propose the incorporation of multimodal distillation into the student RGB model using teacher models. Each teacher model is trained independently on source and target data for its respective modality. Leveraging only unlabeled target data during multimodal distillation enhances the student model’s adaptability to the target domain. We further introduce ensemble masked inference, a technique that reduces the number of input tokens through masking. In this approach, ensemble prediction mitigates the performance degradation caused by masking, effectively addressing the second issue. Our approach outperformed the state-of-the-art CD-FSL approaches with a substantial margin on multiple egocentric datasets, improving by an average of 6.12/6.10 points for 1-shot/5-shot settings while achieving times faster inference speed. Project page: https://masashi-hatano.github.io/MM-CDFSL/
@@
Despite these advancements, a scarcity of egocentric datasets persists in various domains, such as industry and the medical field. Cross-domain approaches [54, 2, 29] effectively transfer the knowledge acquired from the training on large-scale datasets (i.e., source domain) to the target domain. Since the actions taken in the target domain may differ from those in the source domain, the time-consuming manual annotation of the action labels is required. Therefore, cross-domain few-shot learning (CD-FSL) [27, 26, 16, 67, 66] emerges as a crucial strategy, merging cross-domain adaptability with the efficiency of few-shot learning [47, 55, 7, 25, 49, 65] which only uses a few labeled samples on the target domain. Recently, using unlabeled data on the target domain has proven effective in enhancing the adaptability against the domain without increasing the annotation cost [37, 19, 46]. These approaches further improve the task performance in the CD-FSL setting.
@@
Inference Cost. Despite the advancements in the CD-FSL task expanding for videos [46], processing densely-sampled input frames with temporal-aware operations is necessary, making the process computationally expensive [24, 57, 58]. This computational intensity hampers the practical applications on edge devices with limited resources. Therefore, reducing the inference time is essential for egocentric action recognition.
@@ 2 Related Work
Cross-Domain Few-Shot Learning. Guo et al. [14] introduced the Cross-Domain Few-Shot Learning (CD-FSL) challenge, which assesses the capability of deep neural networks to adapt image classification to novel domains that present classes not encountered during their training. Research has explored numerous strategies, such as fine-tuning [4], feature-wise transformation [53], prototype learning [59], and the training with multiple domains [28], all aimed at enhancing the networks’ generalization capabilities. Among these strategies, using unlabeled target data stands out for its ability to increase adaptability to target domains, offering a practical approach akin to unsupervised domain adaptation [15, 9], as the annotation for the target data is not required.
@@ 3.1 Problem Definition
The goal of the cross-domain few-shot learning (CD-FSL) task with multimodal input and unlabeled target data is to classify novel classes within the target dataset , leveraging both a labeled source dataset and an unlabeled target dataset . Both and comprise data across modalities. Ensure there is no class overlap between the source and target datasets to maintain the integrity of the few-shot learning paradigm. To infer novel classes in the target dataset, the data is split into a support set , which contains data from classes for the -way -shot task, and a query set consisting of data from only the classes of the support set, same as the standard few-shot learning setting.
@@ 3.2 Domain-adapted and Class Discriminative Feature Pretraining
Specifically, given an input consisting frames of image size with a channel size for modality RGB, optical flow, hand pose, respectively, we train a VideoMAE [52], which consists of encoder (feature extractor) and decoder , per modality . Following the training framework of VideoMAE, the tokenized input data are randomly masked with the masking ratio by Tube Masking and reconstructed as follows:
@@ 3.3 Multimodal Distillation
Next, we further improve the adaptability to the target domain and few-shot action recognition accuracy in the cross-domain setting by distilling the domain-adapted multimodal features to the RGB feature extractor . Using multiple modalities helps mitigate the domain shift between source and target domain compared to using only RGB modality, as visual information is susceptible to lighting, background, and appearance variations. Incorporating additional modalities like optical flow and hand pose introduces complementary information that is less sensitive to these visual changes. Furthermore, distilling the multimodal features, which are domain-adapted and class-discriminative, into the RGB modality reduces the model complexity while alleviating the domain gap. It aims to imbue the RGB model with domain-adapted characteristics of multimodal learning without the overhead of processing and integrating multiple modality data types during inference.
@@ 4.1 Datasets
4.2 Implementation Details
Training. In the pretraining stage, we train the model, which consists of ViT-S models, pretrained on the Kinetics-400 dataset [21] and a classifier head, which is attached to the ViT-S backbone, for all modalities for 100 epochs. For training settings, we generally follow the VideoMAE [52]. During the multimodal distillation stage, we train the student RGB model for 100 epochs using the AdamW optimizer [31], with a peak learning rate of , linearly increased for the first epochs of the training and decreased to until the end of training with cosine decay [30]. Note that we linearly scaled the peak learning rate with respect to the overall batch size. Regarding the parameters for the loss function, we empirically adapt the balancing hyperparameters in Eq. 2 to for RGB, and for optical flow and hand pose input modality.
@@ 4.3 Comparison Methods
Random Initialization is used as our baseline for the experiment. This method entails learning a logistic regression classifier on top of an untrained VideoMAE encoder, providing a fundamental comparison point for evaluating the efficacy of more sophisticated training strategies.
@@
| Method | Runtime (ms) | GFLOPs | Memory (MiB) | 1-shot | 5-shot | ||||||
| EPIC | MEC | WEAR | Avg. | EPIC | MEC | WEAR | Avg. | ||||
| Random Initialization | 22.1 | 68.5 | 2782 | 29.20 | 23.10 | 25.96 | 26.09 | 40.31 | 27.04 | 38.71 | 35.35 |
| STARTUP++ [37] | 35.17 | 26.85 | 39.15 | 33.73 | 50.23 | 34.05 | 59.88 | 48.05 | |||
| Dynamic Distillation++ [19] | 36.96 | 27.87 | 35.84 | 33.56 | 53.78 | 37.87 | 56.23 | 49.29 | |||
| CDFSL-V [46] | 38.17 | 25.94 | 39.10 | 34.40 | 53.72 | 35.79 | 58.28 | 49.26 | |||
| Ours | 9.64 | 37.0 | 968 | 41.97 | 28.34 | 51.25 | 40.52 | 58.70 | 37.80 | 69.57 | 55.36 |
@@ 4.4 Quantitative Comparison
We compare the performance of 5-way 1-shot and 5-shot top-1 action recognition accuracy and inference time with the prior methods, trained on the Ego4D [12] dataset as the source domain, on three egocentric datasets as the target domain: EPIC-Kitchens [5] in the cooking domain, MECCANO [43] in the industrial-like domain, and WEAR [1] in the outdoor workout domain. We report the accuracy and inference speed of our proposed method with the mask ratio of of , using the ensemble number . Tab. 2 shows that the proposed method consistently outperforms the state-of-the-art CD-FSL with unlabeled target data methods regarding action recognition accuracy and inference speed.
Few-Shot Action Recognition Accuracy. Our proposed model outperforms the CDFSL-V by points and the Dynamic Distillation++ by points on the average of three datasets regarding the 5-shot action recognition accuracy. Significant improvement can be seen in 5-shot accuracy on the WEAR dataset (from to ) as our method leverages the multiple modalities during training, further mitigating the domain gap between source and target. Also, our model outperforms the CDFSL-V by points and the Dynamic Distillation++ by points on average of the 1-shot setting, even though only one labeled target sample from the support set is available during the few-shot training stage.
Inference Time. From Tab. 2, our approach achieved x faster inference speed than the previous approaches while achieving state-of-the-art few-shot classification performance. The approaches in Tab. 2 employ the same architecture (ViT-S) and the same number of input tokens; thus, their inference times remain consistent across implementations. Our approach enhances efficiency and retains high accuracy, making it particularly advantageous for applications demanding online processing without compromising prediction accuracy.
@@ Appendix 0.A Datasets
Appendix 0.B Visualization of Feature Representations
@@
Appendix 0.C Loss Weight Ablation
5-shot |
| 1 | 54.58 |
| 0.1 | 56.68 |
| 0.05 | 57.07 |
| 0.01 | 52.40 |

### v3

@@ Abstract
We address a novel cross-domain few-shot learning task (CD-FSL) with multimodal input and unlabeled target data for egocentric action recognition. This paper simultaneously tackles two critical challenges associated with egocentric action recognition in CD-FSL settings: (1) the extreme domain gap in egocentric videos (e.g., daily life vs. industrial domain) and (2) the computational cost for real-world applications. We propose MM-CDFSL, a domain-adaptive and computationally efficient approach designed to enhance adaptability to the target domain and improve inference cost. To address the first challenge, we propose the incorporation of multimodal distillation into the student RGB model using teacher models. Each teacher model is trained independently on source and target data for its respective modality. Leveraging only unlabeled target data during multimodal distillation enhances the student model’s adaptability to the target domain. We further introduce ensemble masked inference, a technique that reduces the number of input tokens through masking. In this approach, ensemble prediction mitigates the performance degradation caused by masking, effectively addressing the second issue. Our approach outperformed the state-of-the-art CD-FSL approaches with a substantial margin on multiple egocentric datasets, improving by an average of 6.12/6.10 points for 1-shot/5-shot settings while achieving times faster inference speed. Project page: https://masashi-hatano.github.io/MM-CDFSL/
@@
Despite these advancements, a scarcity of egocentric datasets persists in various domains, such as industry [43, 48, 44, 49] and the medical field [62, 8]. Cross-domain approaches [57, 2, 29] effectively transfer the knowledge acquired from the training on large-scale datasets (i.e., source domain) to the target domain. Since the actions taken in the target domain may differ from those in the source domain, the time-consuming manual annotation of the action labels is required. Therefore, cross-domain few-shot learning (CD-FSL) [27, 26, 16, 71, 70] emerges as a crucial strategy, merging cross-domain adaptability with the efficiency of few-shot learning [50, 58, 7, 25, 52, 69] which only uses a few labeled samples on the target domain. Recently, using unlabeled data on the target domain has proven effective in enhancing the adaptability against the domain without increasing the annotation cost [37, 19, 47]. These approaches further improve the task performance in the CD-FSL setting.
@@
Inference Cost. Despite the advancements in the CD-FSL task expanding for videos [47], processing densely-sampled input frames with temporal-aware operations is necessary, making the process computationally expensive [24, 60, 61]. This computational intensity hampers the practical applications on edge devices with limited resources. Therefore, reducing the inference cost is essential for egocentric action recognition.
@@ 2 Related Work
Cross-Domain Few-Shot Learning. Guo et al. [14] introduced the Cross-Domain Few-Shot Learning (CD-FSL) challenge, which assesses the capability of deep neural networks to adapt image classification to novel domains that present classes not encountered during their training. Research has explored numerous strategies, such as fine-tuning [4], feature-wise transformation [56], prototype learning [63], and the training with multiple domains [28], all aimed at enhancing the networks’ generalization capabilities. Among these strategies, using unlabeled target data stands out for its ability to increase adaptability to target domains, offering a practical approach akin to unsupervised domain adaptation [15, 10], as the annotation for large-scale target data is not required.
@@ 3.1 Problem Definition
The goal of the cross-domain few-shot learning (CD-FSL) task with multimodal input and unlabeled target data is to classify novel classes within the target dataset , leveraging both a labeled source dataset and an unlabeled target dataset . Both and comprise data across modalities. Ensure there is no class overlap between the source and target datasets to maintain the integrity of the few-shot learning scheme. To infer novel classes in the target dataset, the data is split into a support set , which contains data from classes for the -way -shot task, and a query set consisting of data from only the classes of the support set, same as the standard few-shot learning setting.
@@ 3.2 Domain-adapted and Class Discriminative Feature Pretraining
Specifically, given an input consisting frames of image size with a channel size for modality RGB, optical flow, hand pose heatmap, respectively, we train a VideoMAE [55], which consists of encoder (feature extractor) and decoder , per modality . Following the training framework of VideoMAE, the tokenized input data are randomly masked with the masking ratio by Tube Masking and reconstructed as follows:
@@ 3.3 Multimodal Distillation
Next, we further improve the adaptability to the target domain by distilling the domain-adapted multimodal features to the RGB feature extractor . Using multiple modalities helps mitigate the domain shift between source and target domain compared to using only RGB modality, as visual information is susceptible to lighting, background, and appearance variations. Incorporating additional modalities like optical flow and hand pose introduces complementary information that is less sensitive to these visual changes. Furthermore, distilling the multimodal features, which are domain-adapted and class-discriminative, into the RGB modality reduces the model complexity while alleviating the domain gap. It aims to imbue the RGB model with domain-adapted characteristics of multimodal learning without the overhead of processing and integrating multiple modality data types during inference.
@@ 4 Experiments
In this section, we elaborate datasets used for training and evaluation, comparison methods, quantitative comparison, and ablation study of our proposed design. (See Suppl. for implementation details and more results).
@@ 4.1 Datasets
| Method | 1-shot | 5-shot | ||||
| EPIC | MEC | WEAR | EPIC | MEC | WEAR | |
| Random Initialization | 29.20
.37 |
23.10
.24 |
25.96
.27 |
40.28
.42 |
27.04
.28 |
38.71
.36 |
| VideoMAE [55] | 35.07
.41 |
27.75
.31 |
44.65
.38 |
47.13
.43 |
35.92
.33 |
63.92
.35 |
| STARTUP
35.18
.43 |
26.84
.30 |
39.15
.35 |
50.24
.45 |
34.05
.31 |
59.88
.36 |
| Dynamic Distill
36.96
.43 |
27.87
.30 |
35.84
.32 |
53.78
.47 |
37.87 .33 | 56.23
.35 |
| CDFSL-V [47] | 38.17
.44 |
26.03
.29 |
39.11
.35 |
53.72
.91 |
35.64
.32 |
58.27
.36 |
| Ours | 41.97 .46 | 28.34 .30 | 51.25 .40 | 58.70 .90 | 37.80 .46 | 69.57 .37 |
@@ 4.2 Comparison Methods
Random Initialization is used as our baseline for the experiment. This method entails learning a logistic regression classifier on top of an untrained VideoMAE encoder.
VideoMAE [55] is used as our baseline, whose parameters are initialized with pretrained on the source (Ego4D) and fine-tuned on the support set.
@@ 4.3 Quantitative Comparison
We compare the performance of 5-way 1-shot and 5-shot top-1 action recognition accuracy and inference cost with the prior methods, trained on the Ego4D [12] dataset as the source domain, on three egocentric datasets as the target domain: EPIC-Kitchens [5] in the cooking domain, MECCANO [43] in the industrial-like domain, and WEAR [1] in the outdoor workout domain. We report the accuracy and inference cost of our proposed method with the mask ratio of of , using the ensemble number . Tab. 2 and Tab. 3 show that the proposed method consistently outperforms the state-of-the-art CD-FSL with unlabeled target data methods regarding action recognition accuracy and inference cost.
Few-Shot Action Recognition Accuracy. Our proposed model outperforms the CDFSL-V by points and the Dynamic Distillation++ by points on the average of three datasets regarding the 5-shot action recognition accuracy. Significant improvement can be seen in 5-shot accuracy on the WEAR dataset (from to ) as our method leverages the multiple modalities during training, further mitigating the domain gap between source and target. Also, our model outperforms the CDFSL-V by points and the Dynamic Distillation++ by points on average of three datasets in the 1-shot setting, even though only one labeled target sample from the support set is available during the few-shot training stage.
Inference Cost. As shown in Tab. 3, our approach achieves x faster inference speed (ms) than the previous approaches while achieving state-of-the-art few-shot classification performance. Furthermore, our method reduces 46 of the theoretical computational cost (GFLOPs), and use only 34 of memory consumption (MiB) compared to the previous methods. The existing approaches in Tab. 3 employ the same architecture (ViT-S) and the same number of input tokens; thus, their inference times remain consistent across implementations. Our approach enhances efficiency and retains high accuracy, making it particularly advantageous for applications demanding online processing without compromising prediction accuracy.
@@ 4.4 Ablation Analysis
Adaptability of the Pre-trained Model. Our teacher encoders in the multimodal distillation depend on pre-trained models. The distillation process using target domain data fails if the pre-trained models do not adequately capture the target domain characteristics. We leverage unlabeled target data during pre-training to adapt the model to the target domain. As shown in Tab. 5, ablation study of the target domain data during the pre-training stage verifies that our method prevents the above issue.
@@ 5 Conclusion
Acknowledgements
This work was supported by JST BOOST, Japan Grant Number JPMJBS2409, and Amano Institute of Technology.
@@ Appendix 0.A Datasets
Appendix 0.B Implementation Details
Training. In the pretraining stage, we train the model, which consists of ViT-S models, pretrained on the Kinetics-400 dataset [21] and a classifier head, which is attached to the ViT-S backbone, for all modalities for 100 epochs. For training settings, we generally follow the VideoMAE [55]. During the multimodal distillation stage, we train the student RGB model for 100 epochs using the AdamW optimizer [31], with a peak learning rate of , linearly increased for the first epochs of the training and decreased to until the end of training with cosine decay [30]. Note that we linearly scaled the peak learning rate with respect to the overall batch size. Regarding the parameters for the loss function, we empirically adapt the balancing hyperparameters to for RGB, and for optical flow and hand pose input modality.
Appendix 0.C Visualization of Feature Representations
@@
Appendix 0.D Loss Weight Ablation
| 1 | 0.1 | 0.05 | 0.01 | |
| 5-shot | 54.58 | 56.68 | 57.07 | 52.40 |